The speed of the current in the Mississippi River is 5 mph. A boat travels downstream 26 miles and returns in 10 2/3 hours. What is its speed in still water?
Answer by Paul(988)

10 2/3 hours = 32/3 (total travelling time)
26/(x+5)+26/(x-5)=32/3
26((x+5)+(x-5)=(32((x+5)(x-5)))/3
26(2x)=(32(x^2-25))/3
52x=(32(x^2-25))/3 ---> multiply the whole equation by 3 to remove the fraction.
156x=32(x^2-25)
156x=32x^2-800
32x^2-156x-800

Factor:
(x+25/8) and (x-8)
Remove the negative.

Hence, the speed of the boat in still water is 8mph.
Paul.